# GRINNs: Godunov-Riemann Informed Neural Networks for Learning Hyperbolic Conservation Laws

### v1

@@ 4 Numerical Results
In this section, we assess the performance of GRINNs for the solution of the inverse problem for four benchmark problems of hyperbolic conservation laws. The first is the inviscid Burgers’ equation [9], a benchmark problem in fluid mechanics end beyond (see e.g., [35, 41, 18]). The second problem, the Lighthill-Whitham-Richards (LWR) equation [44, 64], comes from the field of traffic flow, and shares many similarities with the Burgers’ equation. Here, we consider it to demonstrate the use of GRINNs on finding closures when partial information of the unknown flux function are available. For the next two problems, we consider two systems of $D=2$ non-linear conservation laws; the Shallow Water (SW) equations and the Payne-Whitham (PW) model. The SW equations can be derived by “depth-averaging” the Navier-Stokes equations for an incompressible and inviscid fluid under hydrostatic pressure. The SW equations consist one of the easiest systems to step up from scalar to systems of conservation laws. For all the above problems, the Riemann problem can be solved exactly [41, 35]. This is not the case for the last benchmark problem, the PW model [59, 80] of traffic flow, which describes a class of macroscopic second-order traffic models that take into account the individual behavior of the driver [74, 31]. We consider it here to demonstrate that GRINNs can effectively handle hyperbolic systems with source terms.
For training the GRINNs, we followed the procedure described in section 3.3 to collect training/validation/testing data sets from the numerical solutions of the analytically known hyperbolic PDEs. A presentation of the PDEs, and details on data acquisition, are provided in Appendix A. Next, we assessed the numerical accuracy of GRINNs, by comparing the numerical solutions, provided by the high-resolution Godunov-type scheme in section 3.1, for the learned by GRINNs, hyperbolic equation(s) and the analytically known equation(s) in Appendix A.
@@ 4.1 Learning the Burgers’ equation
for any right and left states $q_{r},q_{l}$ in the data set. Then, the loss function in eq. 17 rendering GRINNs conservative, reduces to:
@@
which was used for solving the GRINNs optimization problem in eq. 15.
As already discussed in section 3.3, we considered $L=5$ neurons in the hidden layer and a logistic sigmoid activation function for learning the physical flux function $\mathcal{N}(u)$ with GRINNs. The training results of GRINNs are shown in table 1, where the maximum and mean $\lVert\hat{\mathbf{Q}}_{i}^{n+1}-\mathbf{Q}_{i}^{n+1}\rVert_{1}$ errors and the MSE$(\hat{\mathbf{Q}}_{i}^{n+1},\mathbf{Q}_{i}^{n+1})$ are reported over all $N=100$ cells and $n_{t}=360$/$360$/$1680$ (training/validation/testing) time steps. Clearly, the GRINNs optimization scheme converges, since the maximum $\lVert\hat{\mathbf{Q}}_{i}^{n+1}-\mathbf{Q}_{i}^{n+1}\rVert_{1}$ error is of the order $061E-$ 06.
@@
Table 1: Training, validation and testing errors of the GRINNs optimization problem for all benchmark problems considered. The maximum and absolute $\lVert\hat{\mathbf{Q}}_{i}^{n+1}-\mathbf{Q}_{i}^{n+1}\rVert_{1}$ errors and the mean squared error MSE$(\hat{\mathbf{Q}}_{i}^{n+1},\mathbf{Q}_{i}^{n+1})$ are reported over all cells $i=1,\ldots,N$, time steps $n=1,\ldots,n_{t}$, and state variable dimensions $D$
The numerical accuracy of GRINNs is visualized in fig. 1a,c. Figure 1a displays the absolute errors between the numerical solution, $\hat{u}(t,x)$, of eq. 19 with the learned by the GRINNs flux functional $\mathcal{N}(u)$, and the ground truth one, $u(t,x)$, of the Burgers’ equation in eq. A.1; the latter is the numerical solution displayed in fig. A.1a (in Appendix). As shown, the numerical accuracy provided by the GRINNs is very high, even near the shock wave regions where the error is at most $056E-$ 05. In addition, we demonstrate in fig. 1c that the flux functional learned by the GRINNs is in excellent agreement with that of the Burgers’ equation $u^{2}/2$.
@@
(c) Burgers’ equation: $u^{2}/2$ vs GRINNs $\mathcal{N}(u)$
(d) LWR equation: $v(\rho)$ vs GRINNs $\mathcal{N}(\rho)$
Figure 1: Numerical accuracy of GRINNs for the Burgers’ and LWR equations. (a,b) Absolute errors of the numerical solution provided by the GRINNs learned equations in eqs. 19 and 21 ($\hat{u}$ and $\hat{\rho}$) vs the one provided by Burgers’ and LWR equations in eqs. A.1 and A.2 ($u$ and $\rho$). (c,d) Analytically known flux function and velocity closure vs the $\mathcal{N}(u)$/$\mathcal{N}(\rho)$ functional learned by GRINNs.
@@ 4.2 Learning the velocity closure of the LWR equation
for any right and left states $\rho_{r},\rho_{l}$ provided by the data set. Then, the loss function in eq. 17 rendering GRINNs conservative, reduces to:
@@
which was used for solving the GRINNs optimization problem in eq. 15; similarly to the Burgers’ equation, $L=5$ neurons in the hidden layer and a logistic sigmoid activation function are considered for approximating $\mathcal{N}(\rho)$.
The training results are reported in table 1, again indicating that the GRINNs optimization scheme converges. In particular, the maximum $\lVert\hat{\mathbf{Q}}_{i}^{n+1}-\mathbf{Q}_{i}^{n+1}\rVert_{1}$ error over all $N=100$ cells and $n_{t}=360$/$360$/$1680$ (training/validation/testing) time steps is around $075E-$ 07.
For the assessment of the numerical accuracy of the GRINNs, we compare in fig. 1b the numerical solution, $\hat{\rho}(t,x)$, of the unknown eq. 21 with the learned by the GRINNs flux functional $\rho\mathcal{N}(\rho)$ and the ground true one, $\rho(t,x)$, of the LWR equation as shown in fig. A.1c. The differences are minor, with the absolute error being very low all over the time-space domain, even near the shock wave regions (the maximum there is $051.2E-$ 05). We further demonstrate in fig. 1d that the velocity closure approximated by the GRINNs is almost identical with the linear one of the LWR equation, $v(\rho)=v_{max}(1-\rho)$.
@@ 4.3 Learning the pressure closure of the SW equations
for any right and left states $(h_{r},q_{r})$, $(h_{l},q_{l})$ provided by the data set. Then, the loss function in eq. 17 that renders GRINNs conservative, reduces to:
@@
where, in this benchmark problem, $\bar{\mathbf{Q}}_{i}^{n}$ requires only the computation of $\bar{h}$ for every cell, since $\bar{q}$ is given by eq. 24. We thus used the loss function in eq. 25 to solve the GRINNs optimization problem in eq. 15. The approximation of the physical flux function $\mathcal{N}(h,q)$ shares the same architecture as in previous benchmark problems. The resulting training results are reported in table 1, where the small training/validation/testing errors indicate the convergence of the GRINNs optimization scheme. In fact, the maximum $\lVert\hat{\mathbf{Q}}_{i}^{n+1}-\mathbf{Q}_{i}^{n+1}\rVert_{1}$ error over all $N=200$ cells and $n_{t}=180$/$180$/$840$ (training/validation/testing) time steps for both state variables $D=1,2$ is around $061E-$ 06.
The numerical accuracy of GRINNs is visualized in fig. 2Figure 2a,b display the absolute errors, for each variable of the system, between the numerical solution of eq. 23 with the learned by the GRINNs flux functional $\mathcal{N}(h,q)$ and the ground truth one, as computed from the SW equations in eq. A.4; the latter is the numerical solution displayed in fig. A.2a,b. Again, the GRINNs provide very high numerical accuracy for both $h$ and $q$ variables, even near the shock wave and rarefaction regions where the error is at most $051.5E-$ 05. In addition, the excellent agreement of the analytically known pressure closure, $P(h)$, and the GRINNs learned functional, $\mathcal{N}(h,q)$, is displayed in fig. 2c. In fact, GRINNs are capable to learn the independence of the pressure closure to $q$; $\mathcal{N}(h,q)$ does not change along that $q$-axis.
@@
(c) $P(h)$ vs GRINNs $\mathcal{N}(h,q)$
Figure 2: Numerical accuracy of GRINNs for the SW equations. (a,b) Absolute errors of the numerical solution provided by the GRINNs learned equations in eq. 23 ($\hat{h}$ and $\hat{q}$) vs the one provided by SW equations in eq. A.4 ($h$ and $q$). (c) Analytically known pressure closure $P(h)$ vs the $\mathcal{N}(h,q)$ functional learned with GRINNs.
@@ 4.4 Learning the pressure closure of the PW equations
which also agrees with the one of the known PW equations in eq. A.9. According to Proposition 1, to render GRINNs conservative, we again assume the arithmetic average $\bar{\rho}=(\rho_{r}+\rho_{l})/2$ and require eq. 10 to hold, which implies the same simplified loss function as the one in eq. 25. We then use this loss function to solve the GRINNs optimization problem in eq. 15.
The training results are reported in table 1; see row denoted as “PW, $P(\rho,q)$”. It is clearly shown that the GRINNs optimization problem does not converge in as low values as in the other problems. In fact, the maximum $\lVert\hat{\mathbf{Q}}_{i}^{n+1}-\mathbf{Q}_{i}^{n+1}\rVert_{1}$ error over all $N=100$ cells and $n_{t}=360$/$360$/$4080$ (training/validation/testing) time steps for both state variables $D=1,2$, is not negligible; at the order of $\mathcal{O}(10^{-3})$. In fact, the pressure closure of the PW equations $P(\rho)$ is poorly approximated by the GRINNs with $\mathcal{N}(\rho,q)$; see fig. D.1d. As a result, the numerical solution of eq. 26 is not accurate, as shown in fig. D.1a,b,c, especially near the shock wave where significant numerical errors accumulate.
This discrepancy occurs because the Roe linearized matrix $\hat{\mathbf{A}}_{i-1/2}(\mathbf{q}_{l},\mathbf{q}_{r})$ for the PW equations in eq. A.9 does not result from the Jacobian $\partial_{\mathbf{u}}\mathbf{f}(\bar{\mathbf{q}})$ computed at the average state $\bar{\mathbf{q}}=[\bar{\rho},\bar{q}]^{\top}$; see discussion in Section C.4. In all other benchmark problems, the Roe linearized matrix $\hat{\mathbf{A}}_{i-1/2}(\mathbf{q}_{l},\mathbf{q}_{r})$ has this property and, thus, GRINNs succeed to solve the inverse problem with a high accuracy.
When GRINNs do not converge
Clearly, when no information is available about the hyperbolic system from which the data come from, such an issue can only be captured by the incapability of GRINNs to converge. The latter provides a hint on wrong predictor selection. We thus attempt to find desired pressure closure in eq. 26 as a function of only $\rho$ and not $q$; i.e., $\mathcal{N}(\rho)$.
By considering $\mathcal{N}(\rho)$ in eq. 26, GRINNs are conservative when eqs. 9 and 10 in Proposition 1 are satisfied. Since no change is employed to the known functions, the average state $\bar{q}$ is again provided by eq. 27. However, requiring eq. 10 to hold for the new $\mathbf{f}_{U}([\rho,q]^{\top})=[0,\mathcal{N}(\rho)]^{\top}$, implies:
@@
The latter equality is similar to the one of the PW equation in eq. A.9 and, in fact, allows us to substitute the above expression of $\partial_{\rho}\mathcal{N}(\bar{\rho})$ into the computation of the Roe linearized matrix of GRINNs.
As a result, the RH condition is explicitly satisfied, and thus the GRINNs optimization problem in eq. 15 can be solved without including the loss function term $\mathcal{L}_{RH}$ in eq. 17, and without using the arithmetic average $\bar{\rho}=(\rho_{r}+\rho_{l})/2$. The training results, derived by solving this simplified GRINNs optimization problem, are reported in table 1; see row denoted as “PW, $P(\rho)$”. It is therein clearly shown that considering a density-dependent pressure closure makes the GRINNs optimization problem converge to very small training/validation/testing errors. In fact, all errors are of magnitude, as low as the errors for the other benchmark problems.
The numerical accuracy of GRINNs is visualized in fig. 3.
@@
(c) $P(\rho)$ vs GRINNs $\mathcal{N}(\rho)$
Figure 3: Numerical accuracy of GRINNs, assuming a pressure closure of the form $\mathcal{N}(\rho)$, for the PW equations. (a,b) Absolute errors of the numerical solution provided by the GRINNs learned equations in eq. 26 for $\mathcal{N}(\rho)$ ($\hat{\rho}$ and $\hat{q}$) vs the one provided by PW equations in eq. A.6 ($\rho$ and $q$). (c) Analytically known pressure closure $P(\rho)$ in eq. A.7 vs the $\mathcal{N}(\rho)$ functional learned with GRINNs.
Figure 3a,b display the absolute errors between the numerical solution of eq. 26 with the learned by the GRINNs density-dependent flux functional $\mathcal{N}(\rho)$ and the one computed from the PW equations in eq. A.6; the latter numerical solution is displayed in fig. A.2d,e. Very high numerical accuracy is provided by the GRINNs for both $\rho$ and $q$ variables, even near the shock wave and rarefaction regions. In these regions, the error was 2-3 orders higher when considering GRINNs for a density $\mathcal{N}(\rho,q)$; see fig. D.1a,b. Finally, the excellent agreement of the analytically known non-linear pressure closure in eq. A.7, $P(\rho)$, and the GRINNs learned functional, $\mathcal{N}(\rho)$ is shown in fig. 3c.

### v2

@@ 4 Numerical Results
In this section, we assess the performance of GoRINNs for the solution of the inverse problem for four benchmark problems of hyperbolic conservation laws. The first is the inviscid Burgers’ equation [9], a benchmark problem in fluid mechanics end beyond (see e.g., [35, 41, 18]). The second problem, the Lighthill-Whitham-Richards (LWR) equation [44, 64], comes from the field of traffic flow, and shares many similarities with the Burgers’ equation. Here, we consider it to demonstrate the use of GoRINNs on finding closures when partial information of the unknown flux function are available. For the next two problems, we consider two systems of $D=2$ non-linear conservation laws; the Shallow Water (SW) equations and the Payne-Whitham (PW) model. The SW equations can be derived by “depth-averaging” the Navier-Stokes equations for an incompressible and inviscid fluid under hydrostatic pressure. The SW equations consist one of the easiest systems to step up from scalar to systems of conservation laws. For all the above problems, the Riemann problem can be solved exactly [41, 35]. This is not the case for the last benchmark problem, the PW model [59, 80] of traffic flow, which describes a class of macroscopic second-order traffic models that take into account the individual behavior of the driver [74, 31]. We consider it here to demonstrate that GoRINNs can effectively handle hyperbolic systems with source terms.
For training the GoRINNs, we followed the procedure described in section 3.3 to collect training/validation/testing data sets from the numerical solutions of the analytically known hyperbolic PDEs. A presentation of the PDEs, and details on data acquisition, are provided in Appendix A. Next, we assessed the numerical accuracy of GoRINNs, by comparing the numerical solutions, provided by the high-resolution Godunov-type scheme in section 3.1, for the learned by GoRINNs, hyperbolic equation(s) and the analytically known equation(s) in Appendix A.
@@ 4.1 Learning the Burgers’ equation
for any right and left states $q_{r},q_{l}$ in the data set. Then, the loss function in eq. 17 rendering GoRINNs conservative, reduces to:
@@
which was used for solving the GoRINNs optimization problem in eq. 15.
As already discussed in section 3.3, we considered $L=5$ neurons in the hidden layer and a logistic sigmoid activation function for learning the physical flux function $\mathcal{N}(u)$ with GoRINNs. The training results of GoRINNs are shown in table 1, where the maximum and mean $\lVert\hat{\mathbf{Q}}_{i}^{n+1}-\mathbf{Q}_{i}^{n+1}\rVert_{1}$ errors and the MSE$(\hat{\mathbf{Q}}_{i}^{n+1},\mathbf{Q}_{i}^{n+1})$ are reported over all $N=100$ cells and $n_{t}=360$/$360$/$1680$ (training/validation/testing) time steps. Clearly, the GoRINNs optimization scheme converges, since the maximum $\lVert\hat{\mathbf{Q}}_{i}^{n+1}-\mathbf{Q}_{i}^{n+1}\rVert_{1}$ error is of the order $061E-$ 06.
@@
Table 1: Training, validation and testing errors of the GoRINNs optimization problem for all benchmark problems considered. The maximum and absolute $\lVert\hat{\mathbf{Q}}_{i}^{n+1}-\mathbf{Q}_{i}^{n+1}\rVert_{1}$ errors and the mean squared error MSE$(\hat{\mathbf{Q}}_{i}^{n+1},\mathbf{Q}_{i}^{n+1})$ are reported over all cells $i=1,\ldots,N$, time steps $n=1,\ldots,n_{t}$, and state variable dimensions $D$
The numerical accuracy of GoRINNs is visualized in fig. 1a,c. Figure 1a displays the absolute errors between the numerical solution, $\hat{u}(t,x)$, of eq. 19 with the learned by the GoRINNs flux functional $\mathcal{N}(u)$, and the ground truth one, $u(t,x)$, of the Burgers’ equation in eq. A.1; the latter is the numerical solution displayed in fig. A.1a (in Appendix). As shown, the numerical accuracy provided by the GoRINNs is very high, even near the shock wave regions where the error is at most $056E-$ 05. In addition, we demonstrate in fig. 1c that the flux functional learned by the GoRINNs is in excellent agreement with that of the Burgers’ equation $u^{2}/2$.
@@
(c) Burgers’ equation: $u^{2}/2$ vs GoRINNs $\mathcal{N}(u)$
(d) LWR equation: $v(\rho)$ vs GoRINNs $\mathcal{N}(\rho)$
Figure 1: Numerical accuracy of GoRINNs for the Burgers’ and LWR equations. (a,b) Absolute errors of the numerical solution provided by the GoRINNs learned equations in eqs. 19 and 21 ($\hat{u}$ and $\hat{\rho}$) vs the one provided by Burgers’ and LWR equations in eqs. A.1 and A.2 ($u$ and $\rho$). (c,d) Analytically known flux function and velocity closure vs the $\mathcal{N}(u)$/$\mathcal{N}(\rho)$ functional learned by GoRINNs.
@@ 4.2 Learning the velocity closure of the LWR equation
for any right and left states $\rho_{r},\rho_{l}$ provided by the data set. Then, the loss function in eq. 17 rendering GoRINNs conservative, reduces to:
@@
which was used for solving the GoRINNs optimization problem in eq. 15; similarly to the Burgers’ equation, $L=5$ neurons in the hidden layer and a logistic sigmoid activation function are considered for approximating $\mathcal{N}(\rho)$.
The training results are reported in table 1, again indicating that the GoRINNs optimization scheme converges. In particular, the maximum $\lVert\hat{\mathbf{Q}}_{i}^{n+1}-\mathbf{Q}_{i}^{n+1}\rVert_{1}$ error over all $N=100$ cells and $n_{t}=360$/$360$/$1680$ (training/validation/testing) time steps is around $075E-$ 07.
For the assessment of the numerical accuracy of the GoRINNs, we compare in fig. 1b the numerical solution, $\hat{\rho}(t,x)$, of the unknown eq. 21 with the learned by the GoRINNs flux functional $\rho\mathcal{N}(\rho)$ and the ground true one, $\rho(t,x)$, of the LWR equation as shown in fig. A.1c. The differences are minor, with the absolute error being very low all over the time-space domain, even near the shock wave regions (the maximum there is $051.2E-$ 05). We further demonstrate in fig. 1d that the velocity closure approximated by the GoRINNs is almost identical with the linear one of the LWR equation, $v(\rho)=v_{max}(1-\rho)$.
@@ 4.3 Learning the pressure closure of the SW equations
for any right and left states $(h_{r},q_{r})$, $(h_{l},q_{l})$ provided by the data set. Then, the loss function in eq. 17 that renders GoRINNs conservative, reduces to:
@@
where, in this benchmark problem, $\bar{\mathbf{Q}}_{i}^{n}$ requires only the computation of $\bar{h}$ for every cell, since $\bar{q}$ is given by eq. 24. We thus used the loss function in eq. 25 to solve the GoRINNs optimization problem in eq. 15. The approximation of the physical flux function $\mathcal{N}(h,q)$ shares the same architecture as in previous benchmark problems. The resulting training results are reported in table 1, where the small training/validation/testing errors indicate the convergence of the GoRINNs optimization scheme. In fact, the maximum $\lVert\hat{\mathbf{Q}}_{i}^{n+1}-\mathbf{Q}_{i}^{n+1}\rVert_{1}$ error over all $N=200$ cells and $n_{t}=180$/$180$/$840$ (training/validation/testing) time steps for both state variables $D=1,2$ is around $061E-$ 06.
The numerical accuracy of GoRINNs is visualized in fig. 2Figure 2a,b display the absolute errors, for each variable of the system, between the numerical solution of eq. 23 with the learned by the GoRINNs flux functional $\mathcal{N}(h,q)$ and the ground truth one, as computed from the SW equations in eq. A.4; the latter is the numerical solution displayed in fig. A.2a,b. Again, the GoRINNs provide very high numerical accuracy for both $h$ and $q$ variables, even near the shock wave and rarefaction regions where the error is at most $051.5E-$ 05. In addition, the excellent agreement of the analytically known pressure closure, $P(h)$, and the GoRINNs learned functional, $\mathcal{N}(h,q)$, is displayed in fig. 2c. In fact, GoRINNs are capable to learn the independence of the pressure closure to $q$; $\mathcal{N}(h,q)$ does not change along that $q$-axis.
@@
(c) $P(h)$ vs GoRINNs $\mathcal{N}(h,q)$
Figure 2: Numerical accuracy of GoRINNs for the SW equations. (a,b) Absolute errors of the numerical solution provided by the GoRINNs learned equations in eq. 23 ($\hat{h}$ and $\hat{q}$) vs the one provided by SW equations in eq. A.4 ($h$ and $q$). (c) Analytically known pressure closure $P(h)$ vs the $\mathcal{N}(h,q)$ functional learned with GoRINNs.
@@ 4.4 Learning the pressure closure of the PW equations
which also agrees with the one of the known PW equations in eq. A.9. According to Proposition 1, to render GoRINNs conservative, we again assume the arithmetic average $\bar{\rho}=(\rho_{r}+\rho_{l})/2$ and require eq. 10 to hold, which implies the same simplified loss function as the one in eq. 25. We then use this loss function to solve the GoRINNs optimization problem in eq. 15.
The training results are reported in table 1; see row denoted as “PW, $P(\rho,q)$”. It is clearly shown that the GoRINNs optimization problem does not converge in as low values as in the other problems. In fact, the maximum $\lVert\hat{\mathbf{Q}}_{i}^{n+1}-\mathbf{Q}_{i}^{n+1}\rVert_{1}$ error over all $N=100$ cells and $n_{t}=360$/$360$/$4080$ (training/validation/testing) time steps for both state variables $D=1,2$, is not negligible; at the order of $\mathcal{O}(10^{-3})$. In fact, the pressure closure of the PW equations $P(\rho)$ is poorly approximated by the GoRINNs with $\mathcal{N}(\rho,q)$; see fig. D.1d. As a result, the numerical solution of eq. 26 is not accurate, as shown in fig. D.1a,b,c, especially near the shock wave where significant numerical errors accumulate.
This discrepancy occurs because the Roe linearized matrix $\hat{\mathbf{A}}_{i-1/2}(\mathbf{q}_{l},\mathbf{q}_{r})$ for the PW equations in eq. A.9 does not result from the Jacobian $\partial_{\mathbf{u}}\mathbf{f}(\bar{\mathbf{q}})$ computed at the average state $\bar{\mathbf{q}}=[\bar{\rho},\bar{q}]^{\top}$; see discussion in Section C.4. In all other benchmark problems, the Roe linearized matrix $\hat{\mathbf{A}}_{i-1/2}(\mathbf{q}_{l},\mathbf{q}_{r})$ has this property and, thus, GoRINNs succeed to solve the inverse problem with a high accuracy.
When GoRINNs do not converge
Clearly, when no information is available about the hyperbolic system from which the data come from, such an issue can only be captured by the incapability of GoRINNs to converge. The latter provides a hint on wrong predictor selection. We thus attempt to find desired pressure closure in eq. 26 as a function of only $\rho$ and not $q$; i.e., $\mathcal{N}(\rho)$.
By considering $\mathcal{N}(\rho)$ in eq. 26, GoRINNs are conservative when eqs. 9 and 10 in Proposition 1 are satisfied. Since no change is employed to the known functions, the average state $\bar{q}$ is again provided by eq. 27. However, requiring eq. 10 to hold for the new $\mathbf{f}_{U}([\rho,q]^{\top})=[0,\mathcal{N}(\rho)]^{\top}$, implies:
@@
The latter equality is similar to the one of the PW equation in eq. A.9 and, in fact, allows us to substitute the above expression of $\partial_{\rho}\mathcal{N}(\bar{\rho})$ into the computation of the Roe linearized matrix of GoRINNs.
As a result, the RH condition is explicitly satisfied, and thus the GoRINNs optimization problem in eq. 15 can be solved without including the loss function term $\mathcal{L}_{RH}$ in eq. 17, and without using the arithmetic average $\bar{\rho}=(\rho_{r}+\rho_{l})/2$. The training results, derived by solving this simplified GoRINNs optimization problem, are reported in table 1; see row denoted as “PW, $P(\rho)$”. It is therein clearly shown that considering a density-dependent pressure closure makes the GoRINNs optimization problem converge to very small training/validation/testing errors. In fact, all errors are of magnitude, as low as the errors for the other benchmark problems.
The numerical accuracy of GoRINNs is visualized in fig. 3.
@@
(c) $P(\rho)$ vs GoRINNs $\mathcal{N}(\rho)$
Figure 3: Numerical accuracy of GoRINNs, assuming a pressure closure of the form $\mathcal{N}(\rho)$, for the PW equations. (a,b) Absolute errors of the numerical solution provided by the GoRINNs learned equations in eq. 26 for $\mathcal{N}(\rho)$ ($\hat{\rho}$ and $\hat{q}$) vs the one provided by PW equations in eq. A.6 ($\rho$ and $q$). (c) Analytically known pressure closure $P(\rho)$ in eq. A.7 vs the $\mathcal{N}(\rho)$ functional learned with GoRINNs.
Figure 3a,b display the absolute errors between the numerical solution of eq. 26 with the learned by the GoRINNs density-dependent flux functional $\mathcal{N}(\rho)$ and the one computed from the PW equations in eq. A.6; the latter numerical solution is displayed in fig. A.2d,e. Very high numerical accuracy is provided by the GoRINNs for both $\rho$ and $q$ variables, even near the shock wave and rarefaction regions. In these regions, the error was 2-3 orders higher when considering GoRINNs for a density $\mathcal{N}(\rho,q)$; see fig. D.1a,b. Finally, the excellent agreement of the analytically known non-linear pressure closure in eq. A.7, $P(\rho)$, and the GoRINNs learned functional, $\mathcal{N}(\rho)$ is shown in fig. 3c.